The area of a room is 20 square metre. If the length is increased by 3 metre and the width is increased by 1 metre, the room will double in area. Determine the original dimensions of the room.
Answer by stanbon(48568)

The area of a room is 20 square metre. If the length is increased by 3 metre and the width is increased by 1 metre, the room will double in area. Determine the original dimensions of the room.
-----------------
lw = 20
(l+3)(w+1) = 40
--------------------
lw + l + 3w + 3 = 40
20 + l + 3w + 3 = 40
l + 3w = 17
--------
But l = 20/w
(20/w) + 3w = 17
20 + 3w^2 = 17w
3w^2 - 17w + 20 = 0
w = [17 +- sqrt(17^2 - 4*3*20)]/6
w = [17 +- sqrt(49)]/6
w = [17+-7]/6
w = 4 or w = 5/3
=======
If width = 4, then length = 5
This answer checks out.
----------
If width = 5/3,
lw = 20
l(5/3) = 20
l = 12
This is wrong.
Therefore w = 5/3 is an extraneous answer.
=================
Cheers,
Stan H